# A Multi-Component, Multi-Physics Computational Model for Solving Coupled Cardiac Electromechanics and Vascular Haemodynamics

### v1

@@ Abstract
The circulatory system, comprising the heart and blood vessels, is vital for nutrient transport, waste removal, and homeostasis. Traditional computational models often isolate individual biophysical processes, such as cardiac electromechanics and blood flow dynamics, failing to capture the system’s integrated nature. This paper presents an innovative approach that couples a 3D electromechanical model of the heart with a 3D fluid mechanics model of vascular blood flow. Our file-based partitioned coupling scheme allows these models to run independently while sharing essential data through intermediate files. We validate this approach using two solvers: one for cardiac electromechanics and the other for vascular blood flow. Developed by separate research groups, these solvers emphasise different dynamical scales and utilise distinct discretisation schemes. Numerical simulations using idealised and realistic anatomies show that the implemented coupling scheme is reliable and requires minimal additional computation time relative to advancing individual time steps in the heart and blood flow models. Notably, the coupled model predicts muscle displacement differently than the standalone heart model, highlighting the influence of detailed vascular blood flow on cardiac function. This study presents a paradigm case of how to build virtual human models and digital twins by productive collaboration between teams with complementary expertise.
@@
circulatory system , cardiovascular simulation , partitioned coupling scheme , high-fidelity simulation , high-performance computing , digital twin
@@ 1 Introduction
In this paper, we present an innovative approach to couple a 3D electromechanical model of the heart with a 3D fluid mechanics model of the blood vessels (see Figure 1). We adopt a multi-component coupling strategy rather than rewriting everything into a new, monolithic code, leveraging existing specialised solvers to build a more comprehensive model. To address this complexity, we employ a file-based partitioned coupling scheme, which couples the two models by alternately writing and reading data through intermediate files. This approach allows each model to run independently while sharing essential simulation data, thereby maintaining optimal computational performance.
We establish the validity of this approach by using Alya [41, 5] to model cardiac electromechanics and HemeLB [42, 43] to model vascular blood flow. These models have been developed by separate research groups, emphasise different dynamical scales, and utilise distinct discretisation schemes. By leveraging the excellent scaling performance of these models on supercomputers, this study aims to construct an efficient multi-component, multi-physics model of the cardiovascular system. Moreover, it illustrates how a virtual human [44], also referred to as a human digital twin, can be developed by integrating efforts from different research groups working in complementary, specialised fields.
The subsequent sections of this paper are structured as follows. In Section 2, we present the individual and coupled models of cardiac function and blood flow, along with the simulation tests designed to verify the coupled model. In Section 3, we analyse the results of these tests, focusing on the computational aspects of the coupled model, and discuss the differences between the solutions obtained from the coupled model and those from the individual models. In Section 4, we discuss the limitations of the coupled model. Finally, we conclude our findings in Section 5.
@@ 2.1 Electromechanical Model of the Heart in Alya
To model the ventricular electrical activity and mechanical contraction, we use a fully-coupled electromechanical model developed in Alya [41, 5]: a simulation code designed for solving multi-physics problems in engineering applications using finite element method. The code is designed for high-performance computing and has demonstrated scalability up to 100,000 cores [41].
@@
Let $\Omega^{\mathcal{A}}_{0}\in\mathbb{R}^{3}$ be the end-diastolic configuration of a left ventricle at time $t=0$, also representing our reference configuration, and $T$ the final time. As we sketch in Figure 2, $\Omega^{\mathcal{A}}_{0}$ is bounded by $\partial\Omega^{\mathcal{A}}_{0}=\Gamma^{\mathcal{A}}_{0_{\mathrm{endo}}}\cup\Gamma^{\mathcal{A}}_{0_{\mathrm{epi}}}\cup\Gamma^{\mathcal{A}}_{0_{\mathrm{base}}}$, which are the endocardium, epicardium, and base respectively; we denote by $\bm{n}_{0}$ the outward-pointing normal to the boundary.
@@
$\bm{D}=\frac{\bm{\sigma}}{C_{\mathrm{m}}S_{\mathrm{v}}},=\frac{1}{C_{\mathrm{m}}S_{\mathrm{v}}}\left(\sigma_{f}\tilde{\bm{f}}_{0}\otimes\tilde{\bm{f}}_{0}+\sigma_{s}\tilde{\bm{s}}_{0}\otimes\tilde{\bm{s}}_{0}+\sigma_{n}\tilde{\bm{n}}_{0}\otimes\tilde{\bm{n}}_{0}\right),$
where $\sigma_{f}$, $\sigma_{s}$, and $\sigma_{n}$ are the conductivity in the fibre, sheet and normal directions and $\tilde{\bm{f}}_{0}$, $\tilde{\bm{s}}_{0}$, and $\tilde{\bm{n}}_{0}$ the unit vector field expressing their directions. Cardiac fibres are modelled with Laplace-Dirichlet rule-based methods [47, 48], as depicted in Figure 2. $C_{\mathrm{m}}$ is the membrane capacitance per unit area and $S_{\mathrm{v}}$ the surface-to-volume ratio. $I_{\mathrm{app}}(\bm{X},t)$ is the applied current; $\bm{W}$ and $\bm{C}$ are the right-hand sides of the system of ODEs corresponding to $\bm{w}$ and $\bm{c}$, respectively. The term $I_{\mathrm{ion}}$ collects the ionic currents. For the cellular ionic model, we use the one introduced by O’Hara et al. [49] and incorporate the modifications proposed by Passini et al. [50] and the modified conductance as described by Dutta et al. [51].
Let $\bm{\Phi}:\Omega^{\mathcal{A}}_{0}\times(0,T)\to\mathbb{R}^{3}$ be a deformation operator which maps each point $\bm{X}$ of the reference configuration $\Omega^{\mathcal{A}}_{0}$ onto the current configuration $\Omega^{\mathcal{A}}_{t}=\bm{\Phi}(\Omega^{\mathcal{A}}_{0},t)$, for all $t\in(0,T)$, so that each point of the current configuration is $\bm{x}(\bm{X},t)=\bm{\Phi}(\bm{X},t)$. We define the displacement field $\bm{d}:\Omega^{\mathcal{A}}_{0}\times(0,T)\to\mathbb{R}^{3}$ as $\bm{d}(\bm{X},t)=\bm{x}(\bm{X},t)-\bm{X}$, the deformation tensor $F(\bm{X},t)=\partialderivative{\bm{\Phi}}{\bm{X}}$, and $S(\bm{d},t)$ is the second Piola-Kirchhoff stress tensor. $J$ is the determinant of $F$: $J=\det(F)>0$. The mechanical deformation is modelled by Cauchy’s first equation of motion equipped with initial and boundary conditions:
@@
where $\rho_{0}^{\mathcal{A}}$ is the tissue density. We consider an active stress formulation, for which we can decompose $S$ into an active and a passive term. The excitation-contraction coupling is modelled using the active force generation model introduced by Land et al. [52]. For the passive part, we consider the Holzapfel and Ogden constitutive law [53].
We set a null initial condition in terms of both displacement and stress, as expressed in equation 8. In equation 5, we set Robin-type boundary conditions on the epicardium with $k_{\mathrm{epi}}$ the corresponding normal spring stiffness. We orient the ventricular geometry such that the normal to the base coincides with the $z$ axis. On the base, we set the displacement to zero in the $x$ and $y$ direction, leaving a stress-free condition in the remaining direction $z$. In equation 6, $\hat{\bm{e}}_{x}$ and $\hat{\bm{e}}_{y}$ are the unit vectors in the $x$ and $y$ directions, respectively. On the endocardium, as expressed in equation 7, we set a Neumann boundary condition with the time-dependent ventricular blood pressure $p_{\mathrm{endo}}(t)$. To model the latter, we subdivide the heartbeat into four phases: isovolumetric contraction, ejection, isovolumetric relaxation, and filling (see Figure 2). After an initialization phase that allows the cardiac muscle to be brought to the time of end-diastole, we model each phase of the heartbeat as follows [54]:
@@
    where $V$ is the ventricular volume; $c_{1}$ and $c_{2}$ are two coefficients that allow us to keep the condition of constant volume. The second term, acting on the second-order volume derivative, stabilizes potential spurious oscillations in the pressure [54].
@@
    $\frac{\mathrm{d}p_{\mathrm{endo}}(t)}{\mathrm{d}t}=-r_{1}\frac{\mathrm{d}V(t)}{\mathrm{d}t}-r_{2}\frac{\mathrm{d}^{2}V(t)}{\mathrm{d}t^{2}},$
    where $r_{1}$ and $r_{2}$ are two coefficients that allow us to keep the volume constant.
@@
Spatial discretisation is carried out with linear finite elements. We use a first-order Yanenko operator splitting scheme and the implicit Euler time advancing scheme for electrophysiology (see [5] for additional details). We solve the non-linear mechanical problem with the Newton-Raphson scheme. Time discretisation of the mechanical problem is carried out with the Newmark scheme using the Rayleigh-damping method to damp potential spurious oscillations [57, 5].
@@ 2.2 Fluid Mechanics Model of Vascular Blood Flow in HemeLB
The 3D blood flow in vessels is simulated using HemeLB [42, 58], an open-source fluid flow solver based on the lattice Boltzmann method (LBM) [59, 60]. This solver was verified and validated in previous works examining vascular flows [61, 62, 63]. HemeLB has been developed into versions optimised for CPU and GPU architectures, Both versions have demonstrated excellent scaling characteristics on high-performance computers: the CPU version has utilised over 300,000 CPU cores, while the GPU version has leveraged over 18,000 GPUs [43]. In this study, we use the CPU version of HemeLB [64].
@@
with a dynamic viscosity $\nu_{0}$ given by
@@
Let $\Omega^{\mathcal{H}}\in\mathbb{R}^{3}$ be the fluid domain of a vascular model. It is bounded by the vessel wall $\Gamma_{\mathrm{wall}}^{\mathcal{H}}$, an inlet $\Gamma_{\mathrm{in}}^{\mathcal{H}}$, and a group of outlets $\Gamma_{\mathrm{out}}^{\mathcal{H}}=\{{\Gamma_{\mathrm{out}}^{\mathcal{H}}}_{,i}\}$. As an illustration, the arterial model used in this study is shown in Figure 3.
In the fluid domain, $\Omega^{\mathcal{H}}$, the blood is assumed to be Newtonian with a constant dynamic viscosity ($\nu_{0}$) of 0.0035 Pa s [67] and a constant density ($\rho_{0}^{\mathcal{H}}$) of 1050 kg m-3 [68]. The vessel walls, $\Gamma_{\mathrm{wall}}^{\mathcal{H}}$, are assumed to be rigid by imposing the Bouzidi-Firdaouss-Lallemand boundary condition [69] on the walls, $\Gamma_{\mathrm{wall}}^{\mathcal{H}}$.
@@
At the outlet sections, $\Gamma_{\mathrm{out}}^{\mathcal{H}}$, the Nash pressure boundary condition is imposed [61]. The pressure $P$ is calculated from the flow rate $Q$ using the WK2 model [55, 56], where the flow rate is obtained by integrating the flow velocity in the normal direction over the outlet plane. Using a semi-explicit scheme to discretise the WK2 model equation [72], the pressure at time step $n+1$ is given by
$P^{[n+1]}=\frac{R^{\mathcal{H}}C^{\mathcal{H}}}{R^{\mathcal{H}}C^{\mathcal{H}}+\Delta t^{\mathcal{H}}}P^{[n]}+\frac{R^{\mathcal{H}}\Delta t^{\mathcal{H}}}{R^{\mathcal{H}}C^{\mathcal{H}}+\Delta t^{\mathcal{H}}}Q^{[n]},$ (19)
where $R^{\mathcal{H}}$ and $C^{\mathcal{H}}$ represent the resistance and capacitance of peripheral vessels respectively. The choice of these parameters is different in the simulations for different tests and is therefore provided individually in Sections 2.5.2 and 2.5.3.
Figure 4: Viscosity $\nu$ within a viscous sponge layer as a function of the distance from the outlet centre $r$, normalised by the base dynamic viscosity $\nu_{0}$ and the radius of the sponge layer $R_{v}$ respectively. The amplification factor of the viscosity $\nu/\nu_{0}$ smoothly increases from 1 on the sponge layer surface ($r=R_{v}$) to 1000 at the outlet centre ($r=0$).
We include viscous sponge layers around the outlets to prevent reflections from the boundaries [73]. Each outlet is surrounded by a spherical sponge layer with a radius $R_{v}$ at its midpoint. Within the sponge layer ($r\leq R_{v}$), the dynamic viscosity $\nu$ gradually increases from the base value $\nu_{0}$ at the spherical surface ($r=R_{v}$) to the maximum value $\nu_{\max}$ at the centre ($r=0$). This variation is described by a function of the distance from the outlet centre $r$ as

### v2

@@ Abstract
The circulatory system, comprising the heart and blood vessels, is vital for nutrient transport, waste removal, and homeostasis. Traditional computational models often treat cardiac electromechanics and blood flow dynamics separately, overlooking the integrated nature of the system. This paper presents an innovative approach that couples a 3D electromechanical model of the heart with a 3D fluid mechanics model of vascular blood flow. Using a file-based partitioned coupling scheme, these models run independently while sharing essential data through intermediate files. We validate this approach using solvers developed by separate research groups, each targeting disparate dynamical scales employing distinct discretisation schemes, and implemented in different programming languages. Numerical simulations using idealised and realistic anatomies show that the coupling scheme is reliable and requires minimal additional computation time relative to advancing individual time steps in the heart and blood flow models. Notably, the coupled model predicts muscle displacement and aortic wall shear stress differently than the standalone models, highlighting the importance of coupling between cardiac and vascular dynamics in cardiovascular simulations. Moreover, we demonstrate the model’s potential for medical applications by simulating the effects of myocardial scarring on downstream vascular flow. This study presents a paradigm case of how to build virtual human models and digital twins by productive collaboration between teams with complementary expertise.
@@
circulatory system , cardiovascular simulation , partitioned coupling scheme , myocardial scarring , high-fidelity simulation , high-performance computing , digital twin
@@ 1 Introduction
In this paper, we present an innovative approach to couple a 3D electromechanical model of the heart with a 3D fluid mechanics model of vascular blood flow (see Figure 1). We adopt a multi-component coupling strategy rather than rewriting everything into a new, monolithic code, leveraging existing specialised solvers to build a more comprehensive model. To address this complexity, we employ a file-based partitioned coupling scheme, which couples the two models by alternately writing and reading data through intermediate files. This approach allows each model to run independently while sharing essential simulation data, thereby maintaining optimal computational performance.
We establish the validity of this approach by using Alya [41, 5] to model cardiac electromechanics and HemeLB [42, 43] to model vascular blood flow. These models have been developed by separate research groups, emphasise disparate dynamical scales, utilise distinct discretisation schemes, and are written in different programming languages. By leveraging the excellent scaling performance of these models on supercomputers, this study aims to construct an efficient multi-component, multi-physics model of the cardiovascular system and investigate its potential for simulating realistic physiological and pathological conditions. Moreover, it illustrates how a virtual human [44], also referred to as a human digital twin, can be developed by integrating efforts from different research groups working in complementary, specialised fields.
The subsequent sections of this paper are structured as follows. In Section 2, we present the individual and coupled models of cardiac function and blood flow, along with the simulation tests designed to verify the coupled model. In Section 3, we analyse the results of four simulation tests designed to evaluate different aspects of the coupled model, including verification of the coupling scheme, quantification of coupling effects, and application to studying the impact of myocardial scarring on vascular flow. In Section 4, we discuss the limitations of the coupled model. Finally, we conclude our findings in Section 5.
@@ 2.1 Electromechanical Model of the Heart in Alya
To model the ventricular electrical activity and mechanical contraction, we use a fully-coupled electromechanical model developed in Alya [41, 5]: a simulation code designed for solving multi-physics problems in engineering applications using the finite element method. Written in Fortran, the code is designed for high-performance computing and has demonstrated scalability up to 100,000 cores [41].
@@
Let $\Omega^{\mathcal{A}}_{0}\subset\mathbb{R}^{3}$ be the end-diastolic configuration of a left ventricle at time $t=0$, also representing our reference configuration, and $T$ the final time. As we sketch in Figure 2, $\Omega^{\mathcal{A}}_{0}$ is bounded by $\partial\Omega^{\mathcal{A}}_{0}=\Gamma^{\mathcal{A}}_{0_{\mathrm{endo}}}\cup\Gamma^{\mathcal{A}}_{0_{\mathrm{epi}}}\cup\Gamma^{\mathcal{A}}_{0_{\mathrm{base}}}$, which are the endocardium, epicardium, and base respectively; we denote by $\bm{n}_{0}$ the outward-pointing normal to the boundary.
@@
$\bm{D}=\frac{\bm{\sigma}}{C_{\mathrm{m}}S_{\mathrm{v}}}=\frac{1}{C_{\mathrm{m}}S_{\mathrm{v}}}\left(\sigma_{f}\tilde{\bm{f}}_{0}\otimes\tilde{\bm{f}}_{0}+\sigma_{s}\tilde{\bm{s}}_{0}\otimes\tilde{\bm{s}}_{0}+\sigma_{n}\tilde{\bm{n}}_{0}\otimes\tilde{\bm{n}}_{0}\right),$
where $\sigma_{f}$, $\sigma_{s}$, and $\sigma_{n}$ are the conductivity in the fibre, sheet and normal directions and $\tilde{\bm{f}}_{0}$, $\tilde{\bm{s}}_{0}$, and $\tilde{\bm{n}}_{0}$ the unit vector field expressing their directions. Cardiac fibres are modelled with Laplace-Dirichlet rule-based methods [47, 48], as depicted in Figure 2. $C_{\mathrm{m}}$ is the membrane capacitance per unit area and $S_{\mathrm{v}}$ the surface-to-volume ratio. $I_{\mathrm{app}}(\bm{X},t)$ is the applied current; $\bm{W}$ and $\bm{C}$ are the right-hand sides of the system of ODEs corresponding to $\bm{w}$ and $\bm{c}$, respectively. The term $I_{\mathrm{ion}}$ collects the ionic currents. We use the cellular ionic model proposed by O’Hara et al. [49] and incorporate the modifications proposed by Passini et al. [50] and the modified conductance as described by Dutta et al. [51].
Let $\bm{\Phi}:\Omega^{\mathcal{A}}_{0}\times(0,T)\to\mathbb{R}^{3}$ be a deformation operator which maps each point $\bm{X}$ of the reference configuration $\Omega^{\mathcal{A}}_{0}$ onto the current configuration $\Omega^{\mathcal{A}}_{t}=\bm{\Phi}(\Omega^{\mathcal{A}}_{0},t)$, for all $t\in(0,T)$, so that each point of the current configuration is $\bm{x}(\bm{X},t)=\bm{\Phi}(\bm{X},t)$. We define the displacement field $\bm{d}:\Omega^{\mathcal{A}}_{0}\times(0,T)\to\mathbb{R}^{3}$ as $\bm{d}(\bm{X},t)=\bm{x}(\bm{X},t)-\bm{X}$, the deformation tensor $F(\bm{X},t)=\partialderivative{\bm{\Phi}}{\bm{X}}$, and $J$ is the determinant of $F$: $J=\det(F)>0$. We consider an active stress formulation, for which we can decompose the second Piola-Kirchhoff stress tensor $S$ into a passive and an active term:
$S(\bm{d},T_{\mathrm{act}}(\bm{q}))=S_{\mathrm{pass}}(\bm{d})+S_{\mathrm{act}}(T_{\mathrm{act}}(\bm{q})),$
where $T_{\mathrm{act}}$ is the active tension and $\bm{q}$ a vector of state variables associated with the contractile mechanism in cardiomyocytes [52, 53]. The mechanical deformation is modelled by Cauchy’s first equation of motion equipped with initial and boundary conditions:
@@
[…]
where $\rho_{0}^{\mathcal{A}}$ is the tissue density in the reference configuration. To model $S_{\mathrm{act}}$, i.e. to model the excitation-contraction coupling, we use the active force generation model introduced by Land et al. [53]. This model consists of a set of ODEs describing the local dynamics of the vector $\bm{q}$:
$\frac{\mathrm{d}\bm{q}}{\mathrm{d}t}=\bm{Q}(\bm{q},\bm{c},\lambda_{\mathrm{f}},\dot{\lambda_{\mathrm{f}}})\quad\text{ in }\Omega^{\mathcal{A}}_{0}\times(0,T),$
where $\lambda_{\mathrm{f}}$ is the stretch in the direction of the fibres, $\bm{Q}$ is the right-hand side of the set of ODEs and $T_{\mathrm{act}}=T_{\mathrm{act}}(\bm{q},\lambda_{\mathrm{f}})$. For additional details on the excitation-contraction model, we refer the readers to Land et al. [53] and Levrero-Florencio et al. [52] for its specific implementation in Alya.
To model the passive part $S_{\mathrm{pass}}$, we consider the Holzapfel and Ogden constitutive law [54].
We set a null initial condition in terms of both displacement and stress, as expressed in equation 9. In equation 6, we set Robin-type boundary conditions on the epicardium with $k_{\mathrm{epi}}$ the corresponding normal spring stiffness. We orient the ventricular geometry such that the normal to the base coincides with the $z$ axis. On the base, we set the displacement to zero in the $x$ and $y$ direction, leaving a stress-free condition in the remaining direction $z$. In equation 7, $\hat{\bm{e}}_{x}$ and $\hat{\bm{e}}_{y}$ are the unit vectors in the $x$ and $y$ directions, respectively. On the endocardium, as expressed in equation 8, we set a Neumann boundary condition with the time-dependent ventricular blood pressure $p_{\mathrm{endo}}(t)$.
To model the ventricular blood pressure $p_{\mathrm{endo}}(t)$, we subdivide the heartbeat into four phases: isovolumetric contraction, ejection, isovolumetric relaxation, and filling (see Figure 2). We denote by $V$ be the volume of the blood contained within the left ventricular cavity. After an initialization phase that allows the cardiac muscle to be brought to the time of end-diastole, we model each phase of the heartbeat as follows [52]:
@@
    where $c_{1}$ and $c_{2}$ are two coefficients that allow us to keep the condition of constant volume. The second term, acting on the second-order volume derivative, stabilizes potential spurious oscillations in the pressure [52]. The coefficients $c_{1}$ and $c_{2}$ are selected a posteriori to ensure that these two conditions are met.
@@
    $\frac{\mathrm{d}p_{\mathrm{endo}}(t)}{\mathrm{d}t}=-r_{1}\frac{\mathrm{d}V(t)}{\mathrm{d}t}-r_{2}\frac{\mathrm{d}^{2}V(t)}{\mathrm{d}t^{2}}.$
    In analogy with the isovolumetric relaxation phase, $r_{1}$ and $r_{2}$ are two coefficients that allow us to keep the volume constant and to avoid spurious pressure oscillation, respectively. Both parameters are selected a posteriori to ensure that both conditions are properly met.
@@
In Alya, the volume $V$ of the blood contained within a cavity can be computed using two different methods. The first method applies the divergence theorem, as explained in Levrero-Florencio et al. [52], while the second method relies on the mixed-product approach. In the latter, the open surface $\Gamma^{\mathcal{A}}_{0_{\mathrm{endo}}}$ is closed with a plane $\Gamma^{\mathcal{A}}_{0_{\mathrm{plane}}}$, which is then triangulated. This process forms a closed surface $\Gamma^{\mathcal{A}}_{0_{\mathrm{pool}}}=\Gamma^{\mathcal{A}}_{0_{\mathrm{endo}}}\cup\Gamma^{\mathcal{A}}_{0_{\mathrm{plane}}}$, with $\Gamma^{\mathcal{A}}_{0_{\mathrm{endo}}}\cap\Gamma^{\mathcal{A}}_{0_{\mathrm{plane}}}=\emptyset$. Let $N_{\mathrm{e}}$ be the total number of elements (2D triangles) in the closed surface $\Gamma^{\mathcal{A}}_{0_{\mathrm{pool}}}$, and let $\bm{a}_{i}$, $\bm{b}_{i}$ and $\bm{c}_{i}$ be the position vectors in the current configuration (dependent on $\bm{d}$) for each vertex of each triangle, with $i=1,\dots,N_{\mathrm{e}}$. The volume enclosed by $\Gamma^{\mathcal{A}}_{0_{\mathrm{pool}}}$ is then given by
$V(\bm{d})=\frac{1}{6}\sum_{i=1}^{N_{\mathrm{e}}}\bm{a}_{i}(\bm{d})\cdot(\bm{b}_{i}(\bm{d})\times\bm{c}_{i}(\bm{d})).$ (10)
Notice that this volume depends on the displacement $\bm{d}$. According to our experience, this method provides a faster computation of the volume compared to the one using the divergence theorem.
Spatial discretisation of the partial differential equations is carried out with linear finite elements. We use a first-order Yanenko operator splitting scheme and the implicit Euler time advancing scheme for electrophysiology (see Santiago et al. [5] for additional details). We solve the non-linear mechanical problem with the Newton-Raphson scheme. Time discretisation of the mechanical problem is carried out with the Newmark scheme using the Rayleigh-damping method to damp potential spurious oscillations [57, 5].
@@ 2.2 Fluid Mechanics Model of Vascular Blood Flow in HemeLB
The 3D blood flow in vessels is simulated using HemeLB [42, 58], an open-source fluid flow solver based on the lattice Boltzmann method (LBM) [59, 60]. This solver was verified and validated in previous works examining vascular flows [61, 62, 63]. HemeLB has been developed into versions optimised for CPU and GPU architectures. Written in C++, both versions have demonstrated excellent scaling characteristics on high-performance computers: the CPU version has utilised over 300,000 CPU cores, while the GPU version has leveraged over 18,000 GPUs [43]. In this study, we use the CPU version of HemeLB [64].
@@
with a kinematic viscosity $\nu_{0}$ given by
@@
Let $\Omega^{\mathcal{H}}\subset\mathbb{R}^{3}$ be the fluid domain of a vascular model. It is bounded by the vessel wall $\Gamma_{\mathrm{wall}}^{\mathcal{H}}$, an inlet $\Gamma_{\mathrm{in}}^{\mathcal{H}}$, and a group of outlets $\Gamma_{\mathrm{out}}^{\mathcal{H}}=\{{\Gamma_{\mathrm{out}}^{\mathcal{H}}}_{,i}\}$. As an illustration, the arterial model used in this study is shown in Figure 3.
In the fluid domain, $\Omega^{\mathcal{H}}$, the blood is assumed to be Newtonian with a constant dynamic viscosity of 0.0035 Pa s [67] and a constant density ($\rho_{0}^{\mathcal{H}}$) of 1050 kg m-3 [68]. This results in a kinematic viscosity ($\nu_{0}$) of $3.33\cdot 10^{-6}$ m2/s. The vessel walls, $\Gamma_{\mathrm{wall}}^{\mathcal{H}}$, are assumed to be rigid by imposing the Bouzidi-Firdaouss-Lallemand boundary condition [69] on the walls, $\Gamma_{\mathrm{wall}}^{\mathcal{H}}$.
@@
At the outlet sections, $\Gamma_{\mathrm{out}}^{\mathcal{H}}$, the pressure boundary condition outlined in Nash et al. [61] is imposed. The pressure $P$ is calculated from the flow rate $Q$ using a three-element Windkessel (WK3) model [55, 56, 72], where the flow rate is obtained by integrating the flow velocity in the normal direction over the outlet plane. The WK3 model extends the WK2 model outlined in Lo et al. [70] by introducing an additional resistance in series with the WK2 components. This extra resistance, known as the characteristic impedance $R_{c}$, accounts for the flow resistance of the proximal vessels and improves the accuracy of the captured flow waveforms, particularly in the medium-to-high-frequency range. In addition, spurious oscillations can be avoided by matching the characteristic impedance with the flow resistance of the proximal vessels [73]. The governing equation of the WK3 model is
$R_{p}C_{p}\frac{\mathrm{d}P}{\mathrm{d}t}+P=R_{c}R_{p}C_{p}\frac{\mathrm{d}Q}{\mathrm{d}t}+(R_{c}+R_{p})Q,$ (21)
where $R_{p}$ and $C_{p}$ are components of the WK2 model representing the resistance and capacitance of peripheral vessels, respectively. Following the approach of Grinberg and Karniadakis [74], we discretise equation 21 using a semi-explicit scheme with first-order approximations. The pressure at time step $n+1$ is given by
$P^{[n+1]}=\frac{R_{p}C_{p}}{R_{p}C_{p}+\Delta t^{\mathcal{H}}}P^{[n]}+\frac{(R_{c}+R_{p})\Delta t^{\mathcal{H}}}{R_{p}C_{p}+\Delta t^{\mathcal{H}}}Q^{[n]}+\frac{R_{c}R_{p}C_{p}}{R_{p}C_{p}+\Delta t^{\mathcal{H}}}\Big{(}Q^{[n]}-Q^{[n-1]}\Big{)}.$ (22)
The choice of the parameters $\{R_{c},R_{p},C_{p}\}$ requires problem-specific tuning. The values used for our tests are provided individually in Sections 2.5.2 and 2.5.3.
Figure 4: Viscosity $\nu$ within a viscous sponge layer as a function of the distance from the outlet centre $r$, normalised by the base kinematic viscosity $\nu_{0}$ and the radius of the sponge layer $R_{v}$ respectively. The amplification factor of the viscosity $\nu/\nu_{0}$ smoothly increases from 1 on the sponge layer surface ($r=R_{v}$) to 1000 at the outlet centre ($r=0$).
We include viscous sponge layers around the outlets to prevent reflections from the boundaries [75]. Each outlet is surrounded by a spherical sponge layer with a radius $R_{v}$ at its midpoint. Within the sponge layer ($r\leq R_{v}$), the kinematic viscosity $\nu$ gradually increases from the base value $\nu_{0}$ at the spherical surface ($r=R_{v}$) to the maximum value $\nu_{\max}$ at the centre ($r=0$). This variation is described by a function of the distance from the outlet centre $r$ as